OpenStudy (anonymous):

Gallium is composed of two naturally occurring isotopes: Ga−69 {\rm Ga}-69 (60.108%) and Ga−71 {\rm Ga}-71. The ratio of the masses of the two isotopes is 1.0290. What is the mass of Ga−69 {\rm Ga}-69?

OpenStudy (aaronq):

the mass of an element on the periodic table is the weighted average of all isotopes (in this case 2). averaged mass of Ga= (%Ga-71)*(mass Ga-71)+(%Ga-69)*(mass Ga-69)

OpenStudy (anonymous):

69.723 = x(68.9256) + (1 - x)(70.9247) 69.723 = 68.9256x + 70.9247 - 70.9247x 1.9991x = 1.2017 x = 0.601 or 60.1% Ga-69 1 - x = 0.399 or 39.9% Ga-71 please give medal

OpenStudy (aaronq):

did you even read the question @s.m.madni ? you straight up copied that from here: http://answers.yahoo.com/question/index?qid=20100916153504AA7D6To and it's wrong because it asked for the mass of Ga-69, not their percentages.